# Orientation Matters: Making 3D Generative Models Orientation-Aligned

### v1

@@ 1 Introduction
Humans possess a remarkable ability to imagine the 3D structure of an object (“synthesis”) and infer its properties, such as orientation (“analysis”), from a single image. We intuitively recognize the correct orientation of a car on the road, know how to grasp a cup by its handle, or decide how to place a chair in a room. In cognitive science, this ability is linked to the concept of object constancy—the capacity to mentally reconstruct a canonicalized object despite changes in viewpoint or other variations [25, 15]. This synthesis of an internal object representation enables analysis of its pose, a process often described as analysis by synthesis.
@@
Despite these strides, a critical challenge remains: orientation alignment. Existing methods often neglect this aspect, largely due to inconsistencies in large-scale 3D datasets [6, 56]. As a result, generated models frequently lack standardized canonical orientations—chairs may face different directions, mugs may appear tipped over, and vehicles may be misaligned. This disparity between human perceptual consistency and the output of generative models poses a significant obstacle for downstream tasks such as analysis-by-synthesis, e.g., orientation estimation. To bridge this perceptual gap, it is essential that 3D generative models not only reconstruct object geometry but also generate objects in consistent, semantically meaningful orientations.
@@
In this paper, we introduce a novel task: orientation-aligned 3D object generation—the generation of a 3D object from a single image such that its orientation is consistently aligned both within and across categories, in accordance with common-sense priors. While no existing methods explicitly address this task, a possible workaround involves first generating a 3D object and then applying pose estimation to align its orientation. However, achieving robust and generalizable orientation alignment across diverse object categories remains a significant challenge. Most 3D pose estimation approaches [53, 28, 39] focus on predicting relative poses with respect to predefined 3D CAD models or multi-view reference images. In contrast, our task demands absolute orientation prediction — aligning objects to a canonical coordinate frame that corresponds to intuitive understandings of front, top, and upright directions. While some prior works explore absolute pose estimation [46, 3], they are typically constrained to a limited set of categories, primarily due to the substantial manual effort required to curate orientation-aligned training data. Recent efforts to scale such methods have relied either on labor-intensive human annotations [14, 23] or on the use of Vision-Language Models (VLMs) [51]. However, the former approach remains category-restricted, while the latter is subject to the inherent inaccuracies of VLM-based orientation predictions. Moreover, these strategies necessitate complex post-processing after 3D generation, which not only introduces potential inaccuracies but also reduces user-friendliness and practical applicability.
@@
To address this, we propose to learn canonical 3D object generation by directly fine-tuning pre-trained 3D generative models, thereby avoiding the limitations of such two-stage pipelines. Our key insight is that a sufficiently diverse set of orientation-aligned 3D models can effectively adapt existing generative models from producing arbitrarily oriented outputs to generating objects with consistent, canonical orientations, while maintaining strong generalization capabilities to unseen objects across various categories. To support this, we introduce Objaverse-OA, a new dataset comprising 14,832 3D models spanning 1,008 categories, each aligned to a consistent, common-sense orientation. Leveraging Objaverse-OA, we fine-tune two representative 3D generative models [56, 22] to produce Trellis-OA and Wonder3D-OA, allowing for generating well-aligned 3D objects across a broad spectrum of categories, including those not included in the fine-tuning set.
@@ 2 Related Work
3D Generative Models: Early 3D generation methods [54, 35, 1, 27, 4] typically employed GANs [45] to model 3D distribution, while these methods generate orientation-aligned objects, they are limited to a single category. The recent breakthroughs in 2D diffusion models [5, 11] provide new solutions for 3D generation. Pioneering works DreamFusion [32] and SJC [47] propose to generate 3D models by distilling from a 2D text-to-image generation model. However, these methods and their follow-ups [17, 18, 42, 41, 50] always suffer from low efficiency and multi-face problems due to per-shape optimization and lack of explicit 3D supervision. Recently, methods based on multi-view diffusion models [22, 20, 48, 36, 40, 37, 43, 59] have succeeded in efficiently producing multi-view consistent images via 3D attention. More recently, [24, 56, 63, 61, 58] utilize 3D latent spaces to further improve the geometry quality of the 2D-assisted approaches. However, despite huge progress on quality and efficiency, they all produce 3D models with uncanonical orientations due to orientation misalignment in their 3D training data, like Objaverse [6], Objaverse-XL [7], and TRELLIS-500K [56].
@@
Object Orientation Estimation: One feasible approach to align 3D model orientations is to render the 3D models from a fixed camera and estimate the object orientations in the renderings. Although image-based 3D object pose estimation has been widely researched, most methods [53, 28, 39, 16, 21, 60] focus on predicting relative poses based on known 3D CAD models or reference images. Since the orientation-aligned 3D CAD models and reference images are not available, they cannot estimate the object orientations aligned with common sense. Category-level object pose estimation methods [46, 3, 52, 19, 9, 44] address this problem by generating 3D shape priors from the input images. However, most of them are limited to category-level due to heavy labor effort in constructing orientation-aligned 3D datasets [46, 2, 14, 23]. In contrast, our method can generate orientation-aligned 3D objects across a large number of categories. Concurrently, Orient Anything [51] realizes zero-shot object orientation estimation by automatically constructing large orientation-aligned 3D datasets using advanced Vision Language Models (VLMs). However, it still suffers from generalizability and accuracy due to a lack of training on the object orientation estimation task. Besides, post-processing after 3D generation is costly compared to directly generating orientation-aligned 3D models.
@@ 3 Objaverse-OA Dataset
In this section, we introduce the construction of our dataset, Objaverse-OA. Dataset diversity plays a crucial role in achieving strong generalization capability. To the best of our knowledge, the existing orientation-aligned 3D dataset [23] with the largest category number includes only 200 categories and fewer than 2,000 3D objects. In contrast, our Objaverse-OA dataset contains 14,832 orientation-aligned 3D objects across 1008 categories, which will be made publicly available to the research community. To build this large-scale dataset while maintaining both efficiency and accuracy, we employ a hybrid pipeline that combines Vision-Language Model (VLM) pre-processing with manual correction.
@@
VLM pre-processing: As discovered by Orient Anything [51], advanced VLMs demonstrate the ability to recognize object front views without task-specific training. Since most models in Objaverse primarily vary in the horizontal (yaw) axis, we follow the strategy proposed in Orient Anything: we render each 3D object from four horizontal viewpoints—front, back, left, and right—and use a VLM to identify the correct front view. Based on the identified view, we then rotate the 3D model accordingly to align it to a canonical orientation. Our data processing begins with the Objaverse-LVIS dataset, and we use Gemini-2.0 [31] as the VLM for view recognition. From a total of 46,219 3D models, Gemini successfully identifies front views for 20,664 objects. However, we observe that VLM-based recognition, while promising, still falls short of human-level accuracy due to the absence of fine-tuning and challenges associated with ambiguous or difficult cases. We illustrate the error rate of VLM’s prediction across different categories in Fig. 2. One of the challenging cases is stick-like objects, like spears, keys, and forks, since many of them are not aligned in roll and pitch angles in the Objaverse. Another challenge involves geometrically narrow or thin objects such as fish, bicycles, and water faucets. In such cases, VLM struggles to identify the correct orientation, as it relies solely on visible front-view features without reasoning from side-view context, unlike humans. What’s more, objects with ambiguous front views, like teapots, extinguishers, and cups, can result in inconsistent orientation predictions. To address these issues, we introduce a manual correction step described below.
@@
Manual correction: Starting from the VLM-based alignment, we manually filter or correct the wrong recognition results of the VLM and canonicalize objects with the ambiguous front-view definition. Moreover, to preserve category diversity, we reintroduce objects that were incorrectly filtered out by the VLM, particularly in cases where a category has too few correctly aligned instances. As illustrated in the Figure 2, we manually correct the orientations for about 600 object categories, especially the stick-like objects and objects with ambiguous orientation or unclear front view features. What’s more, we also filter objects with low geometry quality and scenes with multiple objects. Experiments show that the pre-trained 3D generative models can further improve geometry quality after fine-tuning on our dataset.
@@ 4 Orientation-Aligned 3D Object Generation
We choose a state-of-the-art 3D-VAE-based generative model, Trellis [56], as our base model, which can produce fine-grained geometry and appearance aligned with the input image.
Preliminary: As shown in the Figure 3, Trellis [56] adopts three modules for 3D asset generation during inference including sparse structure generator , structured latents generator , and 3D decoder . Given the input image , sparse structure generator produces dense binary 3D grid : , where N is the length of the grid and is the 3D noise sampled from (0,1), which is further converted into active voxels defined as sparse structure. After that, sparse latents generator is used to generate structured latents : , where is the noised sparse structure . Finally, the 3D representation is obtained via 3D decoding: .
Trellis-OA: The original Trellis model [56] is unable to produce orientation-aligned 3D outputs due to the orientation inconsistencies present in its training data. To address this, we fine-tune Trellis using our Objaverse-OA dataset, resulting in Trellis-OA, which generates 3D objects with aligned orientations. Although Trellis comprises several modules for 3D generation, we find that fine-tuning only the sparse structure generator is sufficient for achieving orientation alignment. This is likely because Trellis inherently generates object poses randomly sampled from four orthogonal directions, and our aligned pose distribution resides within this range. As a result, both the pre-trained structured latent generator and the 3D decoder remain compatible with the aligned orientations and do not require additional fine-tuning. Specifically, our fine-tuned sparse structure generator generates canonical sparse structure , which shares aligned orientations. Afterwards, pre-trained structured latents generator generates canonical structured latents and 3D decoder produces final canonical 3D models with orientation aligned. Our experiments demonstrate that we can efficiently fine-tune Trellis to an orientation-aligned one while preserving its 3D priors.
@@ 4.2 Multi-view Diffusion Model
We choose Wonder3D [22] as our base multi-view diffusion model since it is one of the most representative works based on multi-view diffusion frameworks. Note that the recipe for Wonder3D can also be used by other multi-view diffusion based methods.
Preliminary: Multi-view diffusion models are typically fine-tuned from large-scale pre-trained text-to-image models [34] and achieve multi-view consistent via 3D-aware attention mechanisms. Specifically, given an input image , the multi-view diffusion model generates multi-view images , where is a sampled noise vector. Here, the poses of the ground truth of the multi-view images are dependent on the input image , where is set to predict the input image . After that, Wonder3D adopts an optimization method based on NeuS [49] to lift to 3D representation .
Wonder3D-OA: Wonder3D [22] cannot produce orientation-aligned 3D models due to its misaligned training data and input-image related camera settings. To address this, we implement Wonder3D-OA, leveraging our orientation-aligned Objaverse-OA dataset. Given an input image , Wonder3D-OA first generates orientation-aligned, multi-view consistent images , which are then lifted to orientation-aligned 3D representations . Our key change is in the supervision setup: instead of using input-view-dependent ground truth, we render six canonical views (front, front-left, front-right, left, right, back) from fixed camera poses based on Objaverse-OA. This provides consistent orientation references. We adopt LoRA [12] as a lightweight domain adapter to fine-tune the pre-trained Wonder3D, preserving its learned 3D priors from non-canonical data while enabling alignment. Additionally, we make architectural adjustments for improved performance. Originally, Wonder3D aligns the input image with one of the predicted views to inject local features, which fails to work in our canonical camera setting. Inspired by ImageDream [48], we instead employ a pixel injector to integrate local features into the multi-view diffusion model. We further improve efficiency by replacing Wonder3D’s test-time optimization-based 3D lifting module (NeuS) with LGM [43], a recent model tailored for sparse-view 3D reconstruction. Since LGM is trained using 4-views as input, we feed the front, left, right, and back views into LGM for 3D lifting. Finally, as our lifting module performs effectively without the need for normal maps, we omit fine-tuning the cross-domain attention module in Wonder3D, simplifying the pipeline without sacrificing quality.

### v2

@@ 1 Introduction
Humans possess a remarkable ability to imagine the 3D structure of an object (“synthesis”) and infer its properties, such as orientation (“analysis”), from a single image. We intuitively recognize the correct orientation of a car on the road, know how to grasp a cup by its handle, or decide how to place a chair in a room. In cognitive science, this ability is linked to the concept of object constancy—the capacity to mentally reconstruct a canonicalized object despite changes in viewpoint or other variations [Murray2002ShapePR, Lawson1999AchievingVO]. This synthesis of an internal object representation enables analysis of its pose, a process often described as analysis by synthesis.
@@
Despite these strides, a critical challenge remains: orientation alignment. Existing methods often neglect this aspect, largely due to inconsistencies in large-scale 3D datasets [Deitke2022ObjaverseAU, Xiang2024Structured3L]. As a result, generated models frequently lack standardized canonical orientations—chairs may face different directions, mugs may appear tipped over, and vehicles may be misaligned. This disparity between human perceptual consistency and the output of generative models poses a significant obstacle for downstream tasks such as analysis-by-synthesis, e.g., orientation estimation. To bridge this perceptual gap, it is essential that 3D generative models not only reconstruct object geometry but also generate objects in consistent, semantically meaningful orientations.
@@
In this paper, we introduce a novel task: orientation-aligned 3D object generation—the generation of a 3D object from a single image such that its orientation is consistently aligned both within and across categories, in accordance with common-sense priors. While no existing methods explicitly address this task, a possible workaround involves first generating a 3D object and then applying pose estimation to align its orientation. However, achieving robust and generalizable orientation alignment across diverse object categories remains a significant challenge. Most 3D pose estimation approaches [Wen2023FoundationPoseU6, Peng2018PVNetPV, Sun2022OnePoseOO] focus on predicting relative poses with respect to predefined 3D CAD models or multi-view reference images. In contrast, our task demands absolute orientation prediction — aligning objects to a canonical coordinate frame that corresponds to intuitive understandings of front, top, and upright directions. While some prior works explore absolute pose estimation [Wang2019NormalizedOC, Chen2020CategoryLO], they are typically constrained to a limited set of categories, primarily due to the substantial manual effort required to curate orientation-aligned training data. Recent efforts to scale such methods have relied either on labor-intensive human annotations [Krishnan2024OmniNOCSAU, Ma2024ImageNet3DTG] or on the use of Vision-Language Models (VLMs) [Wang2024OrientAL]. However, the former approach remains category-restricted, while the latter is subject to the inherent inaccuracies of VLM-based orientation predictions. Moreover, these strategies necessitate complex post-processing after 3D generation, which not only introduces potential inaccuracies but also reduces user-friendliness and practical applicability.
@@
To address this, we propose to learn canonical 3D object generation by directly fine-tuning pre-trained 3D generative models, thereby avoiding the limitations of such two-stage pipelines. Our key insight is that a sufficiently diverse set of orientation-aligned 3D models can effectively adapt existing generative models from producing arbitrarily oriented outputs to generating objects with consistent, canonical orientations, while maintaining strong generalization capabilities to unseen objects across various categories. To support this, we introduce Objaverse-OA, a new dataset comprising 14,832 3D models spanning 1,008 categories, each aligned to a consistent, common-sense orientation. Leveraging Objaverse-OA, we fine-tune two representative 3D generative models [Xiang2024Structured3L, Long2023Wonder3DSI] to produce Trellis-OA and Wonder3D-OA, allowing for generating well-aligned 3D objects across a broad spectrum of categories, including those not included in the fine-tuning set.
@@ 2 Related Work
3D Generative Models: Early 3D generation methods [Wu2016LearningAP, Schwarz2020GRAFGR, Chan2021EfficientG3, Pavllo2022ShapePA, Chen2024Learning3G] typically employed GANs [Treat2018GENERATIVEAN] to model 3D distribution, while these methods generate orientation-aligned objects, they are limited to a single category. The recent breakthroughs in 2D diffusion models [Croitoru2022DiffusionMI, Ho2020DenoisingDP] provide new solutions for 3D generation. Pioneering works DreamFusion [Poole2022DreamFusionTU] and SJC [Wang2022ScoreJC] propose to generate 3D models by distilling from a 2D text-to-image generation model. However, these methods and their follow-ups [Liang2023LucidDreamerTH, Lin2022Magic3DHT, Tang2023MakeIt3DH3, Tang2023DreamGaussianGG, Wang2023ProlificDreamerHA] always suffer from low efficiency and multi-face problems due to per-shape optimization and lack of explicit 3D supervision. Recently, methods based on multi-view diffusion models [Long2023Wonder3DSI, Liu2023SyncDreamerGM, Wang2023ImageDreamIM, Shi2023Zero123AS, Szymanowicz2023ViewsetD, Shi2023MVDreamMD, Tang2024LGMLM, Xu2024GRMLG] have succeeded in efficiently producing multi-view consistent images via 3D attention. More recently, [Meng2024LT3SDLT, Xiang2024Structured3L, Zhao2025Hunyuan3D2S, Zhang2024CLAYAC, Xu2023DMV3DDM] utilize 3D latent spaces to further improve the geometry quality of the 2D-assisted approaches. However, despite huge progress on quality and efficiency, they all produce 3D models with uncanonical orientations due to orientation misalignment in their 3D training data, like Objaverse [Deitke2022ObjaverseAU], Objaverse-XL [Deitke2023ObjaverseXLAU], and TRELLIS-500K [Xiang2024Structured3L].
@@
Object Orientation Estimation: One feasible approach to align 3D model orientations is to render the 3D models from a fixed camera and estimate the object orientations in the renderings. Although image-based 3D object pose estimation has been widely researched, most methods [Wen2023FoundationPoseU6, Peng2018PVNetPV, Sun2022OnePoseOO, Lee2025Any6DM6, Liu2025HIPPoHI, Yu2025BoxDreamerDB] focus on predicting relative poses based on known 3D CAD models or reference images. Since the orientation-aligned 3D CAD models and reference images are not available, they cannot estimate the object orientations aligned with common sense. Category-level object pose estimation methods [Wang2019NormalizedOC, Chen2020CategoryLO, Wei2023RGBbasedCO, Lin2022CategoryLevel6O, Fan2022ObjectLD, Tian2020ShapePD] address this problem by generating 3D shape priors from the input images. However, most of them are limited to category-level due to heavy labor effort in constructing orientation-aligned 3D datasets [Wang2019NormalizedOC, Chang2015ShapeNetAI, Krishnan2024OmniNOCSAU, Ma2024ImageNet3DTG]. In contrast, our method can generate orientation-aligned 3D objects across a large number of categories. Concurrently, Orient Anything [Wang2024OrientAL] realizes zero-shot object orientation estimation by automatically constructing large orientation-aligned 3D datasets using advanced Vision Language Models (VLMs). However, it still suffers from generalizability and accuracy due to a lack of training on the object orientation estimation task. Besides, post-processing after 3D generation is costly compared to directly generating orientation-aligned 3D models. Concurrently, [Jin2025Oneshot3O] proposes an intra-category object pose canonicalization method and constructs Canonical Objaverse Dataset, which contains 3D objects with canonical poses within categories. However, our work focuses on inter-category object pose canonicalization.
@@ 3 Objaverse-OA Dataset
In this section, we introduce the construction of our dataset, Objaverse-OA. Dataset diversity plays a crucial role in achieving strong generalization capability. To the best of our knowledge, the existing orientation-aligned 3D dataset [Ma2024ImageNet3DTG] with the largest category number includes only 200 categories and fewer than 2,000 3D objects. In contrast, our Objaverse-OA dataset contains 14,832 orientation-aligned 3D objects across 1008 categories, which will be made publicly available to the research community. To build this large-scale dataset while maintaining both efficiency and accuracy, we employ a hybrid pipeline that combines Vision-Language Model (VLM) pre-processing with manual correction.
@@
VLM pre-processing: As discovered by Orient Anything [Wang2024OrientAL], advanced VLMs demonstrate the ability to recognize object front views without task-specific training. Since most models in Objaverse primarily vary in the horizontal (yaw) axis, we follow the strategy proposed in Orient Anything: we render each 3D object from four horizontal viewpoints—front, back, left, and right—and use a VLM to identify the correct front view. Based on the identified view, we then rotate the 3D model accordingly to align it to a canonical orientation. Our data processing begins with the Objaverse-LVIS dataset, and we use Gemini-2.0 [gemini] as the VLM for view recognition. From a total of 46,219 3D models, Gemini successfully identifies front views for 20,664 objects. However, we observe that VLM-based recognition, while promising, still falls short of human-level accuracy due to the absence of fine-tuning and challenges associated with ambiguous or difficult cases. We illustrate the error rate of VLM’s prediction across different categories in Fig. 2. One of the challenging cases is stick-like objects, like spears, keys, and forks, since many of them are not aligned in roll and pitch angles in the Objaverse. Another challenge involves geometrically narrow or thin objects such as fish, bicycles, and water faucets. In such cases, VLM struggles to identify the correct orientation, as it relies solely on visible front-view features without reasoning from side-view context, unlike humans. What’s more, objects with ambiguous front views, like teapots, extinguishers, and cups, can result in inconsistent orientation predictions. To address these issues, we introduce a manual correction step described below.
@@
Manual correction: Starting from the VLM-based alignment, we manually filter or correct the wrong recognition results of the VLM and canonicalize objects with the ambiguous front-view definition. Moreover, to preserve category diversity, we reintroduce objects that were incorrectly filtered out by the VLM, particularly in cases where a category has too few correctly aligned instances. As illustrated in the Figure 2, we manually correct the orientations for about 600 object categories, especially the stick-like objects and objects with ambiguous orientation or unclear front view features. Note that some objects have orientation ambiguity, especially tools like spoons, mugs, and fire extinguishers. During the manual correction process, we refer to the object orientations defined in prior work, specifically ImageNet3D [Ma2024ImageNet3DTG], for the categories it covers. For example, for spoons, which are included in ImageNet3D, we align their poses in our dataset accordingly. For ambiguous objects only included in our dataset, we define canonical poses based on semantic part structures, geometric features, and common knowledge, following the principles established by ImageNet3D and our supplementary material Section A.1. What’s more, we also filter objects with low geometry quality and scenes with multiple objects. Experiments show that the pre-trained 3D generative models can further improve geometry quality after fine-tuning on our dataset.
@@ 4 Orientation-Aligned 3D Object Generation
We choose a state-of-the-art 3D-VAE-based generative model, Trellis [Xiang2024Structured3L], as our base model, which can produce fine-grained geometry and appearance aligned with the input image.
Preliminary: As shown in the Figure 3, Trellis [Xiang2024Structured3L] adopts three modules for 3D asset generation during inference including sparse structure generator , structured latents generator , and 3D decoder . Given the input image , sparse structure generator produces dense binary 3D grid : , where N is the length of the grid and is the 3D noise sampled from (0,1), which is further converted into active voxels defined as sparse structure. After that, sparse latents generator is used to generate structured latents : , where is the noised sparse structure . Finally, the 3D representation is obtained via 3D decoding: .
Trellis-OA: The original Trellis model [Xiang2024Structured3L] is unable to produce orientation-aligned 3D outputs due to the orientation inconsistencies present in its training data. To address this, we fine-tune Trellis using our Objaverse-OA dataset, resulting in Trellis-OA, which generates 3D objects with aligned orientations. Although Trellis comprises several modules for 3D generation, we find that fine-tuning only the sparse structure generator is sufficient for achieving orientation alignment. This is likely because Trellis inherently generates object poses randomly sampled from four orthogonal directions, and our aligned pose distribution resides within this range. As a result, both the pre-trained structured latent generator and the 3D decoder remain compatible with the aligned orientations and do not require additional fine-tuning. Specifically, our fine-tuned sparse structure generator generates canonical sparse structure , which shares aligned orientations. Afterwards, pre-trained structured latents generator generates canonical structured latents and 3D decoder produces final canonical 3D models with orientation aligned. Our experiments demonstrate that we can efficiently fine-tune Trellis to an orientation-aligned one while preserving its 3D priors.
@@ 4.2 Multi-view Diffusion Model
We choose Wonder3D [Long2023Wonder3DSI] as our base multi-view diffusion model since it is one of the most representative works based on multi-view diffusion frameworks. Note that the recipe for Wonder3D can also be used by other multi-view diffusion based methods.
Preliminary: Multi-view diffusion models are typically fine-tuned from large-scale pre-trained text-to-image models [Rombach2021HighResolutionIS] and achieve multi-view consistent via 3D-aware attention mechanisms. Specifically, given an input image , the multi-view diffusion model generates multi-view images , where is a sampled noise vector. Here, the poses of the ground truth of the multi-view images are dependent on the input image , where is set to predict the input image . After that, Wonder3D adopts an optimization method based on NeuS [Wang2021NeuSLN] to lift to 3D representation .
Wonder3D-OA: Wonder3D [Long2023Wonder3DSI] cannot produce orientation-aligned 3D models due to its misaligned training data and input-image related camera settings. To address this, we implement Wonder3D-OA, leveraging our orientation-aligned Objaverse-OA dataset. Given an input image , Wonder3D-OA first generates orientation-aligned, multi-view consistent images , which are then lifted to orientation-aligned 3D representations . Our key change is in the supervision setup: instead of using input-view-dependent ground truth, we render six canonical views (front, front-left, front-right, left, right, back) from fixed camera poses based on Objaverse-OA. This provides consistent orientation references. We adopt LoRA [Hu2021LoRALA] as a lightweight domain adapter to fine-tune the pre-trained Wonder3D, preserving its learned 3D priors from non-canonical data while enabling alignment. Additionally, we make architectural adjustments for improved performance. Originally, Wonder3D aligns the input image with one of the predicted views to inject local features, which fails to work in our canonical camera setting. Inspired by ImageDream [Wang2023ImageDreamIM], we instead employ a pixel injector to integrate local features into the multi-view diffusion model. Specifically, Wonder3D employs a 3D dense self-attention mechanism with a shape of (, 6, c, , ) across six views within a transformer layer, where is the batch size, c is the number of feature channel, and are the image resolution. Our pixel controller modifies this to (, 7, c, , ), incorporating the input image as an additional view. We further improve efficiency by replacing Wonder3D’s test-time optimization-based 3D lifting module (NeuS) with LGM [Tang2024LGMLM], a recent model tailored for sparse-view 3D reconstruction. Since LGM is trained using 4-views as input, we feed the front, left, right, and back views into LGM for 3D lifting. Finally, as our lifting module performs effectively without the need for normal maps, we omit fine-tuning the cross-domain attention module in Wonder3D, simplifying the pipeline without sacrificing quality.